Find the dy/dx and simply the result, if possible



a. y=√x-1/√x


b. y=x²+π²+xπ


c.y=x² sec x


d.y=sinx-1/cosx


e.y=1/ex+2

1
Expert's answer
2022-03-07T11:08:03-0500

a.


"y'=(\sqrt{x}-\dfrac{1}{\sqrt{x}})'=\dfrac{1}{2\sqrt{x}}+\dfrac{1}{2x\sqrt{x}}"

"=\dfrac{x+1}{2x\sqrt{x}}=\dfrac{\sqrt{x}(x+1)}{2x^2}"

b.


"y'=(x^2+\pi^2+x\pi)'=2x+\pi"

c.


"y'=(x^2\sec x)'=2x\sec x+x^2(-\dfrac{1}{\cos ^2 x})(-\sin x)"

"=2x\sec x+x^2\sec x\tan x"

d.


"y'=(\sin x-\dfrac{1}{\cos x})'=\cos x-(-\dfrac{1}{\cos ^2 x})(-\sin x)"

"=\cos x-\dfrac{\tan x}{\cos x}"

e.


"y'=\dfrac{1}{e^x+2}=-\dfrac{e^x}{(e^x+2)^2}"